Drews age is 8 years less than 3 times Ginas age. Drews is 28 years old how old is Gina?
Answer by BrittanyM(80)

If Drew's age is eight years less than three times Gina's age:
3G - 8 = D
And if Drew's age is 28:
3G - 8 = 28
Now we can solve for G, the unknown variable.
Subtract the negetive eight from both sides:
3G -(-8) = 28 -(-8)
3G = 36
Now divide both sides by three:
G = 12
So Gina is twelve years old.